Can someone help me with these, please?
1) squrt y+7=2
2) squrt 5x+5-5=0
Thanks!!!! :)

Answer by sprolden(40)

1) squrt y+7=2
1- Square each side of the equation. (sqrt y+7)^2 = (2)^2
2- You should get y+7 = 4
3- Subtract 7 from both sides y+7-7 = 4-7
4- You should get y = -3
2) squrt 5x+5-5=0
1- First of all, 5-5 is 0, so you're left with sqrt 5x = 0
2- Square both sides (sqrt 5x)^2 = (0)^2
3- You should get 5x = 0
4- Divide both sides by 5 5/5 x = 0/5
5- You should get x = 0